Heidelberg weekly meeting

03.410 (Physikalisches Institut, Heidelberg)


Physikalisches Institut, Heidelberg


We will have the meeting via Zoom: https://cern.zoom.us/j/92591061032?pwd=MkZzY2g1cG1YZGFwK244c210dk44dz09

Note: this is a temporary room until we get the institute account.

We will keep HeiCONF as a backup: https://heiconf.uni-heidelberg.de/sch-nqu-dd2   Access Code: 858824

Vidyo will be used as a back-up. Room extension 10823320 or use this link.


Tentative minute taker list at https://docs.google.com/spreadsheets/d/15oo_a_1WPxPBOtlH0LDEKG1128y1O2lfswpUAvPCQOY/edit#gid=0

    • 14:00 14:45
      Round Table 45m

      Reports of work progress

      Speakers: Adrian Herkert (Physikalisches Institut, Uni Heidelberg) , Alena Weber (Universität Heidelberg) , Andre Schoening (University Heidelberg, Institute of Physiscs) , Annie Meneses Gonzalez (PhD Student) , Benjamin Weinläder (PI Heidelberg) , David Maximilian Immig (Universität Heidelberg) , Dohun Kim (Physikalisches Institut University of Heidelberg) , Frank Meier Aeschbacher (Universität Heidelberg) , Heiko Augustin (PI Uni Heidelberg) , Lars Noehte (PI Heidelberg) , Luigi Vigani (University of Oxford) , Lukas Mandok (Physikalisches Institut Heidelberg) , Marius Menzel, Sebastian Dittmeier (Physikalisches Institut, Universität Heidelberg) , Sebastian Preuß (PI Uni Heidelberg) , Thomas Theodor Rudzki (Physikalisches Institut Heidelberg)

      Presentation by Alena:

      • Topic is "RadHard Hit Counting Chip"
      • 1st results of testbeam
        • carbon at almost highest possible energie
          • can resolve spills
          • xy map shows nice beam spot (good focus)
        • protons (high energy, low intensity
          • broader spot, more to the left
        • carbon (lower energy as before) only columns
          • very uniform in columns
          • for high energy even better
      • projections in columns is working very nicely, good for HIT
      • can check with such a device during treatment if everything is fine

      Round Table:


      • just started to continue working
      • @Benni: are you at the institute, yes today and tomorrow
      • finishing analysis of recorded data


      • end of last year some of the counting chips were given to irradiation
        In Jan, they are expected to come back, so radiation hardness can be studied


      • finished PCB for HDI studies last year, no feedback from workshop yet
      • sent out chips (run2020) to Uk
      • This week, started talk for HighRR (interview)


      • merging all pull request from Hackathon
      • analysing PSI data MP10 (in progress)
        • Andre: when can results can be expected? --> hopefully next week 1st results


      • installed setup for alpha irradiation
      • next week, 1st test expected with Americium source
      • if working, can start with data taking


      • just came back today


      • Ljubljana might ship sensors next week
      • waiting for Optim to tell when sensors arrive, the 625µm that are thinned to 60,70,80 µm (thinning)
      • chips are sent Monday, the earliest


      • still stuck in HD ...


      • on train (not in a tunnel)
      • worked on DAQ
      • DAB handover?
        • Benni: no handover yet, was planned for today, but did not happened
      • has to quarantine untill negative test, test scheduled for Tuesday

      Sebastian D:

      • Hackathon: went well :-)
        • two pull requests merged that solved 5 issues, config works much smoother now
      • workshop will equip FEB with the coils
        • check these devices then in crate
      • reminder: DPG submission deadline for abstracts (15th January)
      • next week: lab safety instruction session will be given. Mail will be sent later. split sessions since 5 people are max. allowed in labs.
      • On 28th January, meeting for students looking for theses. Sebastian will register us.
        • We should fix the date, some volunteers to present the work in our group.
        • for 24th, we need to prepare some kind of a video


      • cooling studies in Brugg:
        • mockup installed: measurements foreseen:
          • 400/200 mW/cm²
          • transient behavior --> turning off/on
          • 1.8/2.2 instead of 2 g/s
          • turning off half ladders
        • 1st measurements next week
      • @Luigi: please order more PCB for spring run


      • Commissiong run: Helium cooling system readiness not clear, technician not available, chance for helium cooling is below 50 % ... chances for air system
      • meeting on spring demonstrator run after next week's meeting


    • 14:45 14:55